What is the cost of debt when bonds incur a $25 charge and have a par value of $1000, a price of $900, a coupon rate of 4.5% and a maturity of 10 years?

Please show formula and work
Answer #1

Bond Par Value, (FV) = $1,000

Bond Price, (PV) = 900 - 25 = $875

Coupon Payment, (PMT) = 0.045(1,000) = $45

Time to Maturity, (N) = 10 years

Cost of Debt = Yield to Maturity,

Calculating Yield to Maturity,

Using TVM calculation,

I = [PV = -875, FV = 1,000, PMT = 45, N = 10]

I = 0.0622

So,

Cost of Debt = 6.22%